please help me to solve this question:the length of a rectangle is twice its breath. if its length is decreased by 5cm and breadth is increased by 5cm, the area of rectangle is increased by 75 square cm. therefore, what is the length of rectangle?
Answer by mananth(16946)

Breadth-----x cm
length-----2x cm
Area = 2x^2
length & breadth increased by 5
breadth = x+5
length = 2x+5
Area = (x+5)(2x-5)
Area increased by 75 cm^2
(x+5)(2x-5)-2x^2= 75
2x^2-5x+10x-25-2x^2=75
5x-25=75
5x=100
x=20
breadth = 20cm
length = 20*2=40 cm